You have been investing $800 a month for the last 8 years. Today, your investment account is worth $95,596. What is your annual rate of return on your investments? Assume monthly compounding.
Answer #1

Here p = $800, i = ? and n = 8*12 = 96

Thus 95,596 = 800*([1+i)^96 - 1)]/i

or 119.50 = [(1+i)^96 - 1]/i

Solving the above we get 1+i = 1.004459. This i = 0.4459%. This is the monthly rate. So the annual rate = 0.4459%*12

= 5.35%